OpenStudy (anonymous):

1) What is the volume, in cubic centimeters, of a cylinder that has a height of 15 cm and a diameter of 12 cm?

OpenStudy (mertsj):

Volume = area of the base times the height. The base of a cylinder is a circle.

OpenStudy (anonymous):

12x15?

OpenStudy (nenadmatematika):

base: \[(12^2 \times \pi)\div4=36\times \pi\] height: 15 volume o cylinder is \[B \times H = 36\times \pi \times15=540\times \pi cm^3\]